# Compton scattering from superstrings

### v1

@@ 3 Electromagnetic Compton amplitudes for all spins
where $\gamma(x,N)\equiv\prod_{m=1}^{N}(-x+m)=(1-x)_{N}$. This amplitude is gauge invariant in the sense that $q_{4,\beta}{\cal A}^{\alpha\beta}_{n}=0=q_{2,\alpha}{\cal A}^{\alpha\beta}_{n}$. The number of the mass level $n$ characterizes the spin of the massive particle under consideration and we can just take $n\to S$ in the above expressions.
There is still one more constraint that we should consider – we require the associated residues to be given by the product of two minimally coupled 3-point amplitudes. For this purpose, let us analyze the s-channel cut (similar considerations apply to the u-channel cut). We take particle 3 to be a small boost of particle 1 (or the inverse) [72]:
@@
where, after using s-channel kinematics and the fact that $q^{2}\lll 1/\alpha^{\prime}$, we find that $\gamma(s\alpha^{\prime},n-1)=\gamma(n,n-1)$ and $\gamma(u\alpha^{\prime},n-1)=\gamma(n-t\alpha^{\prime},n-1)\approx\gamma(n,n-1)$ at leading order. To understand how we should impose consistent factorization as a constraint, let us analyze the corresponding “helicity amplitudes” separately. First let us consider the all plus amplitude. The term in square brackets is just the spin multipole decomposition of the QED Compton amplitude for matter of spin one [22], so this must be proportional to $\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2}/t$ [69]. So the rest of the amplitude should produce something like $m^{2-2S}\langle{\bf 1}{\bf 3}\rangle^{2S-2}$. The only term that furnishes this is given by the contribution $p=j=n-1$:
@@
where we have denoted by a bar such a contribution. We also used s-channel kinematics. Apart from the positive numerical term inside the brackets, this term reproduces the correct s-channel factorization. Now let us focus on the amplitude with opposite helicities. As the term in square brackets in Eq. (18) is roughly the QED Compton amplitude for matter of spin one, this should be proportional to
@@
Let us quote our proposal for the electromagnetic Compton Scattering amplitude with opposite helicities. Denoting by a bar this particular contribution we find, for a massive particle with spin $s=n>0$ integer:
@@
$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$$\equiv$ ${\cal N}n^{2}(2n)^{n-1}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)$ (23)
$\times$ $\sum_{k=0}^{n-1}\frac{(-1)^{n-1}}{\Gamma(k+1)^{2}\Gamma(-k+n)^{2}}\left(\frac{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1-k}\left(\frac{\bigl{(}q_{4}\cdot\epsilon({\bf 1})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{m^{2}}\right)^{k}$
@@
$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}}\epsilon_{a_{s}}({\bf 1})\epsilon_{c_{s}}({\bf 3}).$
Notice that we have added an arbitrary normalization factor ${\cal N}(n)$ (with ${\cal N}(1)=1$) in order to take care of the multiplicative numerical factor that appears in the factorization of the physical channels.
Let us summarize the three constraints we have imposed:

### v2

@@ 3.3 All mass levels
where $\gamma(x,N)\equiv\prod_{m=1}^{N}(-x+m)=(1-x)_{N}$. This amplitude is gauge invariant in the sense that $q_{4,\beta}{\cal A}^{\alpha\beta}_{n}=0=q_{2,\alpha}{\cal A}^{\alpha\beta}_{n}$. The number of the mass level $n$ characterizes the spin of the massive particle under consideration and we can just take $n\to S$ in the above expressions. Moreover, we remark that this prescription is valid to all mass levels $n\geq 1$.
The Compton amplitude given by Eq. (16) is the full quantum-field-theory amplitude after imposing the two aforementioned constraints, that is, messengers with soft momenta and adequate compactification of one of the massive polarization vectors. As we can see, there are many contributions to the full amplitude, many of them are unlikely to play a role in the root-Kerr problem. However, we will now make the following conjecture: The amplitude whose classical limit describes the dynamics of a compact body is secretly hidden in the full quantum amplitude (16) as a specific contribution. To expose such a contribution, it is obvious that the two requirements established above are not enough; instead, there should be at least one extra step.
We now require the associated residues to be given by the product of two minimally coupled 3-point amplitudes. For this purpose, let us analyze the s-channel cut (similar considerations apply to the u-channel cut). We take particle 3 to be a small boost of particle 1 (or the inverse) [146]:
@@
where, after using s-channel kinematics and the fact that $q^{2}\lll 1/\alpha^{\prime}$, we find that $\gamma(s\alpha^{\prime},n-1)=\gamma(n,n-1)$ and $\gamma(u\alpha^{\prime},n-1)=\gamma(n-t\alpha^{\prime},n-1)\approx\gamma(n,n-1)$ at leading order. To understand how we should impose consistent factorization as a constraint, let us analyze the corresponding “helicity amplitudes” separately. First let us consider the all plus amplitude. The term in square brackets is just the spin multipole decomposition of the QED Compton amplitude for matter of spin one [100], so this must be proportional to $\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2}/t$ in the spinor-helicity basis [142]. So the rest of the amplitude should produce something like $m^{2-2S}\langle{\bf 1}{\bf 3}\rangle^{2S-2}$. The only term that furnishes this is given by the contribution $p=j=n-1$:
@@
where we have denoted by a bar such a contribution. We also used s-channel kinematics. Apart from the positive numerical term inside the brackets, this term reproduces the correct s-channel factorization. Now let us focus on the amplitude with opposite helicities. As the term in square brackets in Eq. (20) is roughly the QED Compton amplitude for matter of spin one, in the spinor-helicity basis this should be proportional to
@@
Let us state our proposal for the electromagnetic Compton scattering amplitude. Consider first the amplitude with opposite helicities. Denoting by a bar this particular contribution we find, for a massive particle with spin $s=n>0$ integer:
@@
$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$$\equiv$ ${\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{\Gamma(n)^{2}}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)$ (25)
$\times$ $\left(\frac{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{\bigl{(}q_{4}\cdot\epsilon({\bf 1})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}\right)$
@@
$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}}\epsilon_{a_{s}}({\bf 1})\epsilon_{c_{s}}({\bf 3})$
where we used that
$\sum_{k=0}^{n-1}\frac{B^{k}A^{-k+n-1}}{\Gamma(k+1)^{2}\Gamma(n-k)^{2}}=\frac{A^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{B}{A}\right)}{\Gamma(n)^{2}}$
and $\,{}_{2}F_{1}\left(a,b;c;x\right)$ is the hypergeometric function. Notice that we have added an arbitrary normalization factor ${\cal N}(n)$ (with ${\cal N}(1)=1$) in order to take care of the multiplicative numerical factor that appears in the factorization of the physical channels. That is,
${\cal N}(n)=\left(\frac{4^{n-1}n^{n+1}\Gamma\left(n-\frac{1}{2}\right)[\gamma(n,n-1)]^{2}}{\sqrt{\pi}\Gamma(n)^{3}}\right)^{-1}.$
As for the all plus amplitude, as discussed we obtain the usual result (up to an irrelevant positive constant) which we quote here for completeness:
$\bar{{\cal A}}_{n}({\bf 1},2^{+},{\bf 3},4^{+})=2ie^{2}m^{2-2n}\frac{\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2n}}{(s-m^{2})(u-m^{2})}\left(\frac{n^{2}[\gamma(n,n-1)]^{2}}{\Gamma(n+1)}\right).$ (26)
[…]
In summary, our claim is: Open string-theory amplitudes can potentially contain amplitudes that describe compact astrophysical bodies as specific contributions (among many others) in disguise inside the full amplitude  [^7] Indeed, as argued in Ref. [105], 3-point string amplitudes produce couplings which are “maximally complex” in that it represents the most general form of the three-point amplitude for one massless and two equal mass legs with spin $S$ but with the corresponding couplings assuming definite values. This amplitude includes, of course, the minimal coupling amplitude which describes root-Kerr in the classical limit.. To extract this hidden contribution to the amplitude, all other contributions must be projected out, and we propose that this can be achieved by imposing the following three constraints: